OpenStudy (deadshot):

How do I set this up if I'm solving for x? 3 over x minus 2 over 3 equals 5 over 6

OpenStudy (deadshot):

Is it \[\frac{ 3 }{ x }-\frac{ 2 }{ 3 }=\frac{ 5 }{ 6 }\] or \[\frac{ \frac{ 3 }{ x-2 } }{ 3 }=\frac{ 5 }{ 6 }\]?

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):



OpenStudy (anonymous):